FlyHigh Tech, an early-stage startup, reports to its investors using a balanced scorecard that is prepared at the end of each quarter. During the first and second quarters of the current year, the company had the following results:

Quarter 1 Quarter 2
During the quarter:
Revenue $ 150,000 $ 180,000
Cost of goods sold 100,000 125,000
Sales and marketing costs 50,000 63,000
General and administrative costs 100,000 110,000
Total market revenue 5,000,000 4,500,000
At the end of the quarter:
Cash $ 2,750,000 $ 2,450,000
Number of revenue-generating customers 6,000 7,500
Number of employees 10 15


The following additional information is available:

  • All jobs were filled as of the end of each quarter, and during the second quarter of the year FlyHigh hired seven new employees.
  • Ten percent of customers from quarter one were lost during the second quarter, however the company increased sales and marketing expenditures by over 20 percent, resulting in a net gain in customers for the quarter.


Required:
Compute the following measures that are included in FlyHigh Tech’s second quarter balanced scorecard: (Round "Cash runway" answer to 1 decimal place.)

Required: Compute the following measures that are included in FlyHigh Techs second quarter balanced Scorecard: (Round Cash
Answer #1

Cash burn rate = [ Starting  cash balance - ending  cash balance ] / Number of months

Cash burn rate = [ $ 2,750,000 - $ 2,450,000] / 3 = $ 100,000

Cash runway = Total cash reserve / Burn rate = $ 2,450,000/ $ 100,000 = 24.5 months

Churn rate = customers lost during 2nd quarter / [ Customers at the beginning + new customers during the period ]

Churn rate = 6,000 X 10%/ [ 6,000 + 1,500] = 0.08 or 8% ( i.e 0.08 X 100%)

Churn rate = 6,000 X 10% / [{6,000 - 600 }+ 2,100] = 0.08 or 8%

Market shares = [ firms sales / Total market revenues ] X 100%

Market shares = [$ 180,000/ $ 4,500,000] X 100% = 4%

Customer acquired cost = cost developed to acquire new customers / number of new customers

Customer acquisition cost per customer = [$ 63,000 - $ 50,000 ] / 7,500 - ( 6,000 - 600)

Customer acquisition cost per customer = $ 13,000 / 2,100 = $ 6.19

Employee turn over rate = [Employees who left / average employees ] X 100%

Employees left = Beginning no of employees + new recruited employees - ending no of employees

Employees left = 10 + 7 - 15 = 2

Average employees = [10 + 15] / 2 = 12.5

Employee turn over ratio = [2/ 12.5] X 100% = 16%